If a 87.9 V emf is connected to the terminals A and B and the current in the 4 ? resistor is 13 A, what is the value of the unknown resistor R?

Round your final answer to five significant figures If a 87.9 V emf is connected to the terminals A and B and the current in the 4 Ω resistor is 13 A, what is the value of the unknown resistor R? 7Ω 2Ω 4Ω R= (2
Answer #1

Here ,

for the resistors in parallel

1/Rp = 1/R1 + 1/R2

Rp = R1 * R2/(R1 + R2)

for resistor in series

Rs = R1 + R2 + ..

for the given circuit

R and 2 Ohm are in series

Rs = R + 2

Rs and 7 Ohm are in parallel

(R + 2) * 7/(R + 2 + 7) = Rp

Rp and 4 are in series

Req = 4 + (R + 2) * 7/(R + 2 + 7)

Now , using Ohm's law

Req = E/I

4 + (R + 2) * 7/(R + 2 + 7) = 87.9/13

solving for R

R = 2.56 Ohm

the resistance R is 2.56 Ohm